### v3

# Pz Cats: Photometric redshift catalogs based on DES Y3 BAO sample

Abstract
Over the years, photometric redshift estimation (photo-z) has advanced through various methods. This study evaluates four distinct photo-z estimators—ANNz2, BPZ, ENF, and DNF—using the Dark Energy Survey Y3 BAO Sample. Unlike most studies, we explore selecting optimal galaxies based on their redshift Probability Distribution Function (PDF) by either reducing noise or identifying those approximating a Gaussian distribution. We cross-matched 25,760 galaxies drawn from four spectroscopic surveys with the photo-z sample to comprehend redshift bias and its 68th percentile $\sigma_{68}$. The lowest $\sigma$ for all estimators was found in the range $0.850.79<z_{p}<$ 0.85. Among the estimators, DNF exhibited the greatest bias, while ENF, ANNz2, and BPZ showed decreased precision outside 0.7 to 0.9 redshift range. To select galaxies with minimal bias, ANNz2 emerged as the most reliable algorithm across all criteria. PDFs selection significantly improves colour representation over the spectroscopic sample, underscoring the role of magnitude space in selection. While ANNz2 achieved superior precision, ENF poorly selected Gaussian PDFs, leaving few galaxies for LSS evaluation. Despite smooth PDFs, catastrophic redshift errors were present. Though DNF had the poorest precision, it offered enough galaxies for cosmological use. Subsampling galaxies with secondary peaks less than 30% of the main peak height, termed Small Peaks, showed ANNz2 excelled. The catalogs produced have been published as Pz Cats within the ZENODO repository.
1 Introduction
Redshift estimation can be obtained in two ways via spectroscopy or photometry. Spectroscopic surveys rely on spectrographs and optical fibres, and this requires a long exposure time of a single object to later compare the spectrum to a measured spectrum in a laboratory. Photometric surveys are based on bands/filters in different wavelength ranges. For this second type of survey construction, we need to estimate the redshift of each observed object from the measured magnitudes and compare them to a set of previously observed objects whose spectroscopy and photometry are known.
Photometric surveys are ideal for observing a larger number of objects, rather than spectroscopic ones, and understand the properties of their shape to infer their morphology. The methods used can be template fitting or training-based machine learning methods. Template fitting is based on the inference of photometric redshift (photo-z) that requires a set of known objects that should include all their sub-types based on empirical and/or theoretical models which allow us to predict the photometry from a few parameters, like stellar mass, star formation rate and redshift, among others. Some examples of algorithms are: LePhare [1, 2], BPZ [3], ZEBRA[4], EAZY [5]. Training-based methods require a set of coincident galaxies whose spectroscopic redshifts are known, some examples of algorithms are: TPZ [6], ANNz2 [7], GPZ [8], DNF[9].
Luminous Red Galaxies (LRG) are the best sources for surveys whose goal is to study the Large Scale Structure (LSS). They are correlated with clusters, occupying massive halos [10, 11]. LRGs are bright and red, they have uniform Spectral Energy Distribution (SED) which makes them ideal for photo-z surveys [12]. In this study, we will focus on the Dark Energy Survey (DES) LRGs for the BAO sample. They used the VISTA Hemisphere Survey (VHS) [13] together with DES filters to select the LRGs in the $rzK$-space [14] for redshifts above $0.5$.
There are many thorough photometric redshift verifications made by the DES Collaboration or independent groups like [15, 16, 17]. Our main goal in this paper is to analyse how selecting galaxies through their full redshift Probability Distribution Function (PDF) can affect the effective redshift ($z_{\rm eff}$) of the survey, its precision and sample size. This is different than assuming the PDFs are nearly Gaussian and comparing different photo-z estimators by their redshift variance. Furthermore, we are using a different training spectroscopic sample, with intersections with the ones used by the DES Collaboration, less populated, but still relevant to understand how each estimator can be influenced by the size of the training sample. In terms of precision, we obtain the photo-z errors without assuming any distribution function to the sample’s distribution $N(z)$.
The present study is organized into five sections. Section 2 describes the photometric estimators we used for this analysis. Section 3 describes the error estimation of the photometric methods. Next, we show in section 4 the results for each estimator and their respective cuts, including selecting the best redshift distribution functions and colour cut. The last section 5 summarises the study.
2 Photo-z estimators
In this work, we used the Dark Energy Survey BAO sample of red galaxies from the third year of observations, Dark Energy Survey (DES) Year 3 [18] photometric sample. The dataset collected from the survey comprises a total of $7,081,993$ galaxies, spread over an area of $21,078.79$ square degrees. This amounts to an observed density of approximately $333.60$ galaxies per square degree. The input of each object was used in the three photo-z estimators within a range of $2.00<z<$ 2.0. We test four different algorithms for estimating the Probability Density Function: ANNz2, BPZ, ENF, DNF  [^1] We included the codes used here at: https://github.com/psilvaf/cat_org..
2.1 BPZ
The Bayesian photometric redshift estimation (BPZ) [3] version 1.98b is a Bayesian probability redshift estimation algorithm, a template fitting method. This method is based on the Bayesian theory that all probabilities are conditional. Thus, the redshift of a galaxy given that we have photometric information will be $p(z|\mathbf{m_{0}})$, where $z$ is the redshift we want to find and $m_{0}$ are the magnitudes in each filter.
Applying Bayes’ theorem, we have
$p(z|\mathbf{m_{0}})=\frac{p(\mathbf{m_{0}}|z)p(z)}{p(\mathbf{m_{0}})},$ (2.1)
where $p(\mathbf{m_{0}}|z)$ is the likelihood, the probability that the magnitude data of various filters $\mathbf{m_{0}}$ represent the redshift $z$. The redshift will be the result that maximises the likelihood.
We chose PDFs with $100$ points and the same redshift range as the chosen for the other algorithms.
2.2 ANNz2
It is possible to obtain photo-zs through an Artificial Neural Network (ANN), a training-based method. It forms a map between the input variables and output ones, the connections through response functions are called neurons. ANNz [19] is a software that uses a multilayer perceptron (MLP). Of the many layers of nodes, the inputs that are the magnitudes of $k$ filters, $\bf{m}_{k}$, are in the first layer, while the last ones contain the outputs: the photometric redshift $z_{p}$ and the PDFs.
To obtain $z_{p}$, the data must be divided into training, testing, and evaluation sets. For samples of galaxies without spectroscopic redshift, one needs to match the galaxies to a spectroscopic survey to use for training and testing.
What ANN does is minimise a cost function $E$ that compares the estimated $z_{p}$ according to the weight $\bf{w}$ and the input magnitudes, and the spectroscopic redshift $z_{k}$ of the training set. $E$ is written as follows:
$E=\sum_{k}\left[z_{p}(\bf{w},\bf{m}_{k})-z_{k}\right].$ (2.2)
ANNz2 [7] is an improved version to find the $z_{p}$ PDF. This can be done through a randomised regression that combines machine learning methods (MLM), which are trained and perform better. The MLM with the best performance is chosen as the final estimator.
We used the Random Regression technique, dividing the reference set into four parts, the training with $20,608$ galaxies and three test sets each with $1,000$; $2,000$; and $2,760$ galaxies, respectively. An ensemble of $100$ random MLMs were used in the photo-z estimation, $90$ k-Nearest Neighbours (kNN), and $100$ PDF bins.
2.3 DNF
Another trained-based method is the nearest neighbours method. [9] developed two methods based on the k-nearest neighbour (kNN) algorithm. An Euclidean Neighbourhood Fitting (ENF) treats all the training galaxies as equally important, the estimated redshift will be close to where there are closer galaxies in the magnitude surface. A direction-oriented fitting, on the other hand, considers how the multi-magnitude surface looks depending on its direction, in this case, the direction is the estimated redshift. Imagine a single galaxy with many neighbours with their respective spectroscopic redshift. For a galaxy at the centre of its neighbours, its redshift will point to the direction where the neighbours form a smoother surface.
A good method ensures that the neighbours do not have just similar spectroscopic redshifts, but also a similar multi-magnitude surface. The Directional Neighbourhood Fitting (DNF) [9] provides a better redshift estimator compared to ENF.
The distance between two galaxies $D$ in a magnitude space is:
$D=\sqrt{\sum_{i}^{n}(m_{i}^{t}-m_{i}^{p})^{2}},$ (2.3)
where $m_{i}^{t}$ is the magnitude of the training sample and $m_{i}^{p}$ is the magnitude of the galaxies with redshift to be estimated, $n$ is the number of magnitude bands. This is the distance used with ENF/DNF.
For the DNF estimator, one might consider the angle $\alpha$ formed between two position vectors in magnitude space. The perfect near-neighbour galaxy has $\alpha=0$, where the distance to a neighbour galaxy is $DN=D^{2}\sin^{2}\alpha$. We also chose 100 points to construct the PDFs. For the specifications with ENF/DNF, we chose 100 neighbours and ensured photometric redshifts remained inside the training redshift range.
2.4 Spectroscopic sample
The reference spectroscopic galaxy set consists of $25,760$ galaxies from: the VIMOS Public Extragalactic Redshift Survey (VIPERS) Data Release 2 (PDR-2) (redshift flag Flag2absent2$\geq$ 2≥ 2) [20] W1 and W4 equatorial fields, DEEP2 Galaxy Redshift Survey (Flag Q$>$3) [21], VIMOS VLT Deep Survey (VVDS) [22], and the Sloan Digital Sky Survey (SDSS) eBOSS LRG pCMASS [23] (HIZ_LRG flag [24]) . These galaxies were found by matching their celestial coordinates with a tolerance of $2$ arcseconds. We used the Random Regression technique, dividing the reference set into four parts, the training with $20,608$ galaxies and three test sets each with $1,000$; $2,000$; and $2,760$ galaxies, respectively. An ensemble of $100$ random Machine Learning Methods was used in the photo-z estimation, $90$ kNN, and $100$ PDF bins. In Figure 1, we show the distribution of colours $g-i$ and $r-i$ for the DES observations for its whole BAO sample in black and our matched galaxies in blue, we do not have information for $r-i>1.5$, but this is shown to be irrelevant in the results. Figure 2 shows the redshift distribution of this sample, ideal for an LSS survey with LRG.
Figure 1: Colours $r-i$ vs $g-i$. of the DES galaxies (black) and the matched spec-z sample (blue).
Figure 2: Spec-z sample distribution.
BPZ does not need a training set of spectroscopic redshift, thus the required input information was simply the magnitudes of the galaxies as well as the filters information, we added the DECam filters transmission as a function of wavelength to the BPZ inputs. We chose PDFs with $100$ points and the same redshift range as the chosen for the other algorithms.
We used the same spectroscopic sample described above for DNF/ENF. We also chose 100 points to construct the PDFs. For the specifications with DNF/ENF, we chose 100 neighbours and ensured photometric redshifts remained inside the training redshift range. Because DNF/ENF estimates the redshift according to neighbouring galaxies from the spectroscopic sample, we had to divide the spectroscopic set into $\sim 5000$ training galaxies and the remaining to compute the metrics for these two estimators, otherwise the result would be exactly the spec-z result.
In Table (1), we show the effective redshift $z_{\rm eff}$ defined in [25] and the 68th percentile of the redshift error $\sigma_{z}$, defined as
$\sigma_{z}=\frac{(z_{p}-z)}{(1+z)}.$ (2.4)
The comparison was made with the spectroscopic dataset used from training and testing. For ANNz2, we used the set separated to test for the algorithm pipeline. The BPZ bias was calculated with all the available spectroscopic observations. Lastly, for DNF/ENF, we had to run the estimator with the galaxies of the training set, since DNF/ENF does not have a testing step. We separated 80% for training and 20% for testing DNF/ENF, and the 20% was used to compute the bias between redshifts. DNF/ENF’s $\sigma_{z}^{68}$ matches the results obtained by [17] in the incomplete training set case, which means we are providing sufficient spectroscopic reference galaxies.
Data $z_{\rm eff}$ $\sigma_{z}^{68}$
ANNz2 0.856 ${}^{+0.007}_{-0.021}$
BPZ 0.848 ${}^{+0.009}_{-0.039}$
ENF 0.854 ${}^{+0.018}_{-0.016}$
DNF 0.819 ${}^{+0.054}_{-0.016}$
Table 1: Uncertainty ($\sigma_{z}^{68}$) and effective redshift ($z_{\rm eff}$) of each algorithm using the full survey.
Figure 3: Redshift bias $\langle\sigma\rangle$ w.r.t. averaged redshift $\langle z_{bin}\rangle$ in different bins of each photo-z estimator. ANNz2 (blue line with star symbols), ENF(red line with dots), BPZ (gray line with swuares), and DNF (dotted line). The DNF’s bias is shown in the right y-axis. The error bars were estimated from the 68th percentile of $\langle\sigma\rangle$ distribution around zero.
3 Bias per redshift bin
A good way to compare these estimates is the quantity $\sigma_{68}$, which is the $68\%$ confidence region of the bias $\sigma$:
$\sigma=z_{photo}-z_{train}.$ (3.1)
We plot the bias, Eq. (3.1), per redshift bin in Figure (3) and their respective uncertainty based on the percentile 68 pra the bias distribution, the sample used to compare is the same as the previous plots. The error bars represent well the number of spectroscopic redshifts available per bin; they are smaller for bins with more spec-z. Consequently, the bias and errors are larger for higher redshift, a common lack of available data due to the nature of spectroscopic surveys. The red squares represent the ENF bias per bin, they are larger for $z>0.95$, this is expected once we do not have enough galaxies with these redshift values. The same is true for ANNz2, but the bias is smaller, indicating a better performance of this estimator. DNF is the most biased estimator for most bins. Compared to the DES Collaboration results in [15], we agree that in the bins around $z_{p}=0.8$, the bias increases at $z_{p}>1.0$.
In Table (1), we show the effective redshift $z_{\rm eff}$ described by [25], and the 68th percentile of the redshift error $\sigma_{z}$ 2.4. This first result is based on the full sample. The comparison was made with the spectroscopic data set used in the training and testing. For ANNz2, we used the separated set to test the algorithm pipeline. The BPZ bias was calculated with the available spectroscopic observations. Lastly, for ENF/DNF, we had to run the estimator with the galaxies of the training set, since ENF/DNF does not have a testing step. We separated 80% for training and 20% for testing ENF/DNF, the 20% was used to compute the bias between redshifts.
4 Results
Our results for the four methods are available in the Zenodo repository with DOI: 10.5281/zenodo.14290701. In Figure 4, we show the ANNz2 testing sample compared to the resulting photo-z. The cyan dotted line shows the $\sigma_{99}$ error, the dashed blue line represents the $\sigma_{68}$ error with the expected result. The black and white pixels represent the number of galaxies concentrated with that particular spec-z/photo-z value. The resulting redshift estimation was satisfactory, the sample distribution matches the spectroscopic redshift pattern without much dispersion.
Figure 4: ANNz2 photo-z compared to the spectroscopic training set. The horizontal axis is the spec-z testing sample. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
Figure 5 shows the high bias between the BPZ photo-z and the matched samples $spec-z$. Compared to ANNz2, this estimator has a bigger error, indicated by the wider $\sigma_{68}$ region. The results are consistent with the expected redshift, with little dispersion from the $z_{p}=z_{spec}$ line. For redshifts greater than 1.1, $\sigma_{99}$ deviates from a straight line. The stripped pattern is due to the limitation of the estimator of setting a redshift resolution, we chose a resolution of $0.02$. This same pattern can be seen in other studies using BPZ, like in [26].
Figure 5: BPZ vs spec-z full reference set. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
Paper Algorithm $\sigma_{68}$ Sample Training set size
DES SV [16] ANNz2 (test 1 cut) $\sim$0.049 DES SV 7,249
DES SV [16] BPZ (test 1) $\sim$0.056 DES SV -
DES SV [16] DESDM (test 1) $\sim$0.055 DES SV 7,249
DES Deep Field [17] DNF Complete 0.0373 DESY3 Deep Field 26,883
DES Deep Field [17] DNF Incomplete 0.0500 DESY3 Deep Field 22,651
Table 2: Photo-z estimation using DES objects.
The ENF results are shown in Figure 6. This time the deviation in the errors region happens at $z\sim 0.95$, this indicates that the estimator is not accurate enough for higher redshift values. Finally in Figure 7, DNF seems to overestimate the redshift and has the same issue at higher redshift as ENF. Both algorithms show an asymmetric distribution, the negative $\sigma_{68}$ is closer to the $z_{spec}=z_{p}$ function for the $z<0.9$, these algorithms are overestimating the galaxy’s photo-z. A similar discrepancy appeared in the lower redshift region in [27] for the Sloan Digital Sky Survey photometric catalog.
In table 2, we included results using DES data with different algorithms, from the Scientific Verification and Deep Field. We had to estimate the $\sigma_{68}$ from [16] to the equivalent version of our definition  [^2] Considering their mean spec-z is around 0.75, we divide their $\sigma_{68}$ by (1+z).. Our results show lower $\sigma_{68}$ than most cases. The exception is DNF Complete training (all the available filters from DECAM) set from [17]. Compared to [17]’s DNF incomplete case (only $g,r,i,z$ filters like the BAO sample), we found compatible results.
Figure 6: ENF photo-z vs spec-z. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
Figure 7: DNF photo-z vs spec-z. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
In order to make sure our spectroscopic sample was enough for DNF, we compare, in Figure 8, our result in the x-axis with the BAO sample by the DES Y3 [18] in the y-axis. They are highly correlated for all the redshift values. The DES Collaboration used $\sim 2.2\times 10^{5}$ spectroscopic matched galaxies in their estimation spectra matched [18]. Despite their superior access to spectroscopic data, our analysis is comparable with $\langle\sigma\rangle=\langle z_{Pz\,\,cats}-z_{DESY3}\rangle=-0.003$. Figure 8 shows the 68th percentile for this bias with the same colour scheme used above. The results diverge for $z>0.9$, this means DNF is more sensitive to the training set in this range.
Figure 8: DNF results with our training set vs the DESY3 one. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between the two data sets. The blue dashed-line represents the 68% confidence level in a $z_{DES}-z_{PzCats}$ distribution, and the cyan-dotted one the 99% confidence level.
4.1 Consistency tests based on the redshift Probability Distribution Function
Due to the lack of a bigger number of filters, we may find redshift Probability Distribution Functions (PDF) that are not perfect, such as a Dirac delta function. An accurate but not free-of-systematics PDF ideally looks like a Gaussian distribution, and the realistic case results in a survey with nearly Gaussian and multi-modal PDFs.
Estimator Method No. of galaxies $z_{\rm eff}$ $\sigma_{z}^{68}$ Area (deg2) Galaxies/deg2
ANNz2 Gaussian $2,931,677$ $0.817$ ${}^{+0.014}_{-0.031}$ 21,065.36 139.17
ANNz2 Small peaks $5,133,775$ $0.799$ ${}^{+0.009}_{-0.022}$ 21,065.36 243.71
ENF Gaussian $299,427$ $0.834$ ${}^{+0.040}_{-0.084}$ 21,032.17 14.24
ENF Small peaks $2,256,616$ $0.768$ ${}^{+0.046}_{-0.055}$ 21,065.36 107.12
DNF Gaussian $2,111,461$ 0.868 ${}^{+0.116}_{-0.121}$ 21,078.79 100.17
DNF Small peaks $2,456,190$ $0.761$ ${}^{+0.033}_{-0.142}$ 21,078.79 116.52
BPZ Gaussian $3,973,671$ $0.812$ ${}^{+0.016}_{-0.045}$ 21,065.36 188.64
BPZ Small peaks $5,904,860$ $0.811$ ${}^{+0.008}_{-0.042}$ 21,065.36 280.31
Table 3: All the cuts for each estimator.
Unfortunately, one cannot eliminate all the objects with secondary peaks, but it is possible to exclude the most noisy ones. Here, we computed the number of peaks per PDF and selected galaxies that contain a peak that cannot be larger than 30% of the main peak. Furthermore, we selected PDFs close to a Gaussian distribution. We calculated the statistical moments $\mu_{n}$ using the distributions $PDF(z_{p})$:
[…]
where $n$ is the moment ordinal, we use the second moment to classify the distribution. For a Gaussian, the second moment is the sum of the average value squared and the variance ($\mu^{2}+\sigma^{2}$) [^3] The code is available at: https://github.com/psilvaf/bao_pz.. The mode and the mean are the same in a normal distribution, so we use the $z_{p}$ output of the estimators as $\mu$ and its respective error as $\sigma$. This sample we call Gaussian + "estimator". We shall call Small Peaks the case where the second biggest maximum is not bigger than thirty per cent of the global maximum, this guarantees that the noise is not dominant in the PDF.
Contrasting with the reweighing approach outlined in [28, 29], which constructs the probability density function (PDF) of the sample by using the weights while also avoiding both the need for trimming and the evaluation of the PDFs for individual galaxies, our sample selection methodology is specifically structured to remove unreliable individual PDFs that might exist within the survey dataset. These PDFs are prone to noise due to their under-representation in the magnitude space, leading to potential distortions in data interpretation.
The resulting sub-samples are shown in the table 3. BPZ has the biggest number of selected objects in both criteria. However, because of the Bayesian method, this estimator forces smoother PDFs, but that may not mean it is the best representation of all galaxies, we need to trust that the template available is representative enough of the galaxy’s photometry. ANNz2 shows better precision despite not having a sample as big as BPZ’s. ENF has the worst selection of Gaussian PDFs, with very few galaxies left for an LSS study. Lastly, the worst in precision is DNF but with sufficient galaxies for cosmological analysis. The Small Peaks results are more populated, this time ANNz2 has better precision and more objects than the other algorithms. Choosing the least noisy PDFs may be an advantage to LSS studies for being less selective because the decrease in the number of objects is lower.
The next question is to determine if the selected samples continue to represent robust large-scale structure (LSS) datasets. In Table 3, we have presented both the residual survey area and the density of objects, expressed as the number of galaxies per square degree. It is important to highlight that the spatial footprint of the survey has remained consistent, unaffected by the decrease in the quantity of objects. Nonetheless, there is a crucial caveat regarding the ENF Gaussian sample: the substantial reduction in object count for this particular sample renders it unsuitable for further analyses involving angular power spectra or two-point correlation functions. Our analysis of the galactic distribution through various cuts supports the data in Table 3. The ENF approach shows a decrease in galaxy counts, while an optimal sky direction features increased galaxy density linked to the spec-z sample. In contrast, ANNz2 exhibits the smallest sky distribution post-cuts.
The principal aim of this survey is to identify the BAO signal. However, a reduction in the number of galaxies compromises the signal, and the of this reduction on statistical power in [30]. It is important to highlight that BAO detection can be accomplished using two methodologies: one that operates in three dimensions and another that is transverse to the line of sight. To achieve accurate transverse BAO measurements, it is crucial to have an adequate density of galaxies per square degree. In contrast, for three-dimensional observations, the redshift distribution needs to be robust enough to capture the characteristic feature along the line of sight, irrespective of whether the survey footprint encompasses an extensively filled section of the sky. In particular, the most recent findings from the Dark Energy Spectroscopic Instrument (DESI) [31] provide an exemplary estimation of 3D BAO even when significant gaps are present across all its tracers. In the last column of Table 3, we included the number of galaxies per deg2, the only sample to lose statistics considerably is ENF, all the others are capable of being a BAO sample.
(a) BPZ
(b) ANNz2
(c) ENF
(d) DNF
Figure 9: Comparison of colour distributions for different methods of photo-z (black) estimation and the matched spec-z sample (blue). The contours represent the $68\%$ confidence level of the distribution: (a) BPZ, (b) ANNz2, (c) ENF, and (d) DNF.
(a) BPZ
(b) ANNz2
(c) ENF
(d) DNF
Figure 10: Comparison of colour distributions for different methods of photo-z (black) estimation and the matched spec-z sample (blue). The contours represent the $68\%$ confidence level of the distribution: (a) BPZ, (b) ANNz2, (c) ENF, and (d) DNF.
[…]
From eq. (3.1) 68th percentile, we plot the distribution of $\sigma_{z}$ for all estimators and sample cuts. In Figure 12, we show the distribution of $\sigma_{z}$ for all the samples described here. The black line represents the whole survey, the blue one represents the samples only with Gaussian PDFs, and the grey area includes multimodal PDFs but not too noisy.
(a) ANNZ $\sigma_{z}$
(b) BPZ $\sigma_{z}$
(c) ENF $\sigma_{z}$
(d) DNF $\sigma_{z}$
Figure 12: Comparison of $\sigma_{z}$ values from different models/algorithms. The vertical red-line marks the null difference, the black histogram has all galaxies, gray histogram represents the Small peaks cuts, and the blue histogram the Gaussian one with the corresponding colours of dashed vertical lines representing the 68% confidence level.
The first panel has the ANNz2 normalised bias, Figure 12(a), the vertical lines respecting the distribution colours. The expected normalised bias is a Gaussian distribution with mean zero, the three cases show the centres close to zero, but the distributions are not perfectly Gaussian, and the ±$\pm\sigma_{68}$ are not symmetric. The least accurate sample is the blue (Gaussian) one, the main reason is the loss of statistics. In Figure 12(b), BPZ’s three distributions are very similar and not Gaussian. ENF and DNF (Figure 12(c),12(d)) have the worst normalised bias distribution and biggest ±$\pm\sigma_{68}$.
4.2 Precision with respect to the redshift
Complementary to the results above, we show $\sigma_{68}$ w.r.t. the spectroscopic redshift of the matched galaxies used for testing (all galaxies for BPZ). The result is shown in Figure 13. BPZ in figure 13(a) looses precision when $z>1.0$, for ENF/DNF(figs. 13(c),13(d)) this happens at $z>0.9$. The only photo-z estimator to show consistent and precise results is ANNz2.
(a) BPZ
(b) ANNz2
(c) ENF
(d) DNF
Figure 13: $\sigma^{68}$ vs $z_{spec}$ bin. Blue: positive error. Red: negative error.
This is consistent with Figures 4, 5, 6, and 7. The asymmetry of overestimation by ENF/DNF and underestimation by BPZ, while ANNz2 is nearly uniform for all redshift values.
We know there are certain galaxies that are not well estimated giving catastrophic redshift. Here, we choose $\sigma_{z}>\sigma_{99}$ as a catastrophic result. In table 4 we compare the numbers of galaxies with catastrophic results, ANNz2 has the best performance compared to the other estimators. Then, we compared the PDF performance in the catastrophic results, BPZ distribution modes $z_{mode}$ match exactly the photo-z result, but the other algorithms show discrepancy, with ANNz2 as with the worst result.
Estimator # of galaxies with $\sigma_{z}>\sigma_{99}$ of galaxies with $|z_{mode}-z_{phot}|>0.1$
ANNz2 146 81
BPZ 1,283 0
ENF 1,179 38
DNF 1,084 27
Table 4: Catastrophic results w.r.t the PDF mode for each algorithm.
We compared the relative frequency of galaxies classified with Gaussian PDFs from the testing sample for each algorithm. In Figure 14, we show the number of Gaussians ($N_{Gauss}$) up to $|\sigma_{z}|$ compared to the number of galaxies in the test sample up to $|\sigma_{z}|$. BPZ has nearly 70% Gaussian PDFs with bias bigger than 0.075 and ANNz2 has nearly 60%. ENF reaches a maximum of 25% of Gaussian PDFs for very small bias, after $|\sigma_{z}|>0.025$, the frequency saturates and only 15% of the PDFs are nearly Gaussian. The opposite happens to DNF, it is an improvement compared to ENF, but saturation is below 25%. This result shows an important situation that may occur; even though the PDFs are smooth, there are catastrophic redshift results.
Figure 14: Relative frequency of Gaussian distributions compared to the total number of testing samples w.r.t. the absolute bias ($\sigma_{z}$) for each estimator. BPZ: black, ANNz2: blue, ENF: gray, and DNF: red.
4.3 Filters with respect to the bias
Another important aspect is how $\sigma_{68}$ behaves for each filter. We chose to analyse the behaviour for magnitudes greater than 20. First, we compare the bias normalized by $\sigma_{68}$ in Figure 15. The first panel shows the distribution for the filter $g$, because DNF has the larger error, the distribution (in red) is narrower than the other estimators. ANNz2 and BPZ are centred at zero. DNF and ENF are centred at positive bias. The same pattern is seen in the second top panel for the filter $r$. The filters $i$ and $z$ show a different pattern, ANNz2 is still centred at zero, while BPZ is centred at negative bias and DNF at positive bias. ENF’s filter $i$ is less positive than the other filters. Finally for filter $z$, in the last panel, the only difference from the previous one is BPZ’s distribution which is now more negative.
Figure 15: Redshift ias distribution for each filter normalized by $\sigma^{68}$. Red: DNF, black: BPZ, gray: ENF and blue: ANNz2.
Therefore, the best estimator is ANNz2, its error is consistent for all redshifts. The other estimators are less precise for higher redshift. For LSS analysis this is problematic, the ideal scenario should include a wider range of redshifts.
4.4 Colour cut
In the previous test, we based our $\sigma_{68}$ analysis w.r.t. the colour coverage of our spectroscopic sample, in Figure 1. For each estimator, we cut the redshifts restriction to the colours $41<g-i<$ 4 and $1.40.5<r-i<$ 1.4. Then, we selected the redshifts whose $\sigma_{68}$ is smaller than $10\%$ of the photo-z value. In table 5, there are the $z_{\rm eff}$ considering the colour cut and it is clear that they are compatible with the entire sample photo-z estimation. lastly, the number of objects reduces significantly to DNF, so depending on the use of the catalog, one must understand that an accurate representation of the filters by the training sample may reduce statistics.
Estimator $z_{\rm eff}$ number of galaxies
ANNz2 0.86 7,031,738
BPZ 0.84 7,012,076
ENF 0.82 6,304,106
DNF 0.84 4,727,115
Table 5: Result after the colour cut.
We checked the bias distribution normalised by $\sigma_{68}$ in terms of the faintest part of each filter. This is shown in Figure 15, the only algorithm that is symmetric on the bias is ANNz2, BPZ underestimates the faint galaxies, while DNF and ENF overestimate them.
Fortunately, the criteria we have established for probability density functions have shown to be highly effective in eliminating suboptimal spectroscopic sample representations based on colour. This is evidenced in Figure 9 where we plot the $68\%$ confidence level of color surface distribution, which shows a notable reduction in subrepresented samples once we exclude the PDFs associated with high noise levels. The decrease in such subrepresentation becomes even more pronounced when we apply the Gaussian cut criterion, as depicted in Fig. 10. Within this Gaussian framework, the algorithm ANNz2 emerges as the most accurate among the methods evaluated, whereas ENF experiences a substantial reduction, although this comes at the cost of some statistical accuracy. Ultimately, our careful selection strategy guarantees a reliable representation with respect to colour metrics, ensuring the overall robustness of our analytical approach.
[…]
5 Summary
This study explored the properties of four different photo-z estimators: ANNz2, BPZ, ENF, DNF with the same spectroscopic sample to train and/or test their precision.
We selected 25,760 galaxies from four different spectroscopic surveys and cross-matched them with the DES Y3 BAO sample. These galaxies served to understand the redshift bias $\sigma$ and its 68th percentile $\sigma_{68}$ and are representative in terms of magnitude. We found that within a range of $0.850.79<z_{p}<$ 0.85 there is the lowest $\sigma$ for all the estimators we analysed. DNF has the biggest absolute value of the bias, while ENF, ANNz2 and BPZ lose precision for a redshift range below 0.7 and greater than 0.9.
Computing the statistical metrics of the redshfit errors we found that the smallest bias belong to BPZ and the worst bias belongs to DNF. The least biased bin is $z_{p}\sim 0.8$ for all estimators, while the most biased bins are below 0.7 and above 0.9.
We also investigated the $\sigma_{68}$ metric, a robust metric to estimate the error of the algorithms. ANNz2 and BPZ showed consistent $\sigma_{68}$ for all bins, while ENF and DNF have an asymmetric error distribution, with lower negative error values than positive ones and a considerable error increase for redshift greater than 0.9. As a consistent test, we compared with the DES Collaboration, because our and their results diverge slightly for higher redshift, we conclude that this algorithm is sensitive to the training set available especially in the high redshift end.
In another evaluation we conducted, our objective was to identify the optimal galaxies based on their redshift Probability Density Functions (PDFs). This endeavour resulted in a subset where BPZ identified the highest count of objects. Nevertheless, the Bayesian methodology intrinsic to BPZ tends to produce more homogenised PDFs. Conversely, ANNz2, although possessing a smaller subset than BPZ, demonstrated superior accuracy. On the other hand, ENF generated the least effective selection, associated with Gaussian PDFs, leaving a minimal quantity of galaxies suitable for a Large-Scale Structure (LSS) analysis, only 14 galaxies per degree squared. DNF, while being the least precise of the methods, still retained an adequate number of galaxies for cosmological investigations. Within the Small Peaks sub-sample, ANNz2 outshone the other systems. Given the stringent criteria of the Gaussian subsample, it is less favourable for analyses requiring robust statistical outcomes. A noticeable effect of selecting PDFs is an improved representation in terms of colours compared with the spectroscopic sample; this finding underscores that such selection is significantly influenced by the magnitude space of the sample.
In case one wants to pick the best galaxies by removing the bins with the worst bias, they will find that ANNz2 is the most robust algorithm for all criteria. DNF and ENF have a significant shift from zero. We also showed that even though the PDFs are smooth, there are catastrophic redshift results.
Finally, we conducted an analysis of how bias manifests in the context of faint galaxies. Our findings corroborate the previous analyses, indicating that ANNz2 maintains consistency across all tested redshift ranges, whereas the other estimators exhibit tendencies to either overestimate or underestimate the photometric redshifts. Ultimately, among the four algorithms evaluated, ANNz2 demonstrated success across the majority of the metrics we discussed, with comparatively fewer catastrophic failures. In an upcoming study, we aim to investigate whether this level of success exerts an influence on cosmological outcomes.
6 Code availability
The particular software packages used in this work will be accessible at https://github.com/psilvaf/bao_pz and https://github.com/psilvaf/cat_org.
Acknowledgments
We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E26/210.130/2023.
The authors would like to thank the anonymous referee who provided useful and detailed comments.
PSF thanks Brazilian funding agency CNPq for PhD scholarship GD 140580/2021-2. RRRR thanks CNPq for partial financial support (grant no. $309868/2021-1$).
The authors thank Juan De Vicente (Centro Investigaciones Energéticas, Medioambientales y Tecnológicas) for providing the DNF estimator code.
This project used public archival data from the Dark Energy Survey (DES). Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology FacilitiesCouncil of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, the Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Científico e Tecnológico and the Ministério da Ciência, Tecnologia e Inovação, the Deutsche Forschungsgemeinschaft, and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenössische Technische Hochschule (ETH) Zürich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciències de l’Espai (IEEC/CSIC), the Institut de Física d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universität München and the associated Excellence Cluster Universe, the University of Michigan, the National Optical Astronomy Observatory, the University of Nottingham, The Ohio State University, the OzDES Membership Consortium, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. Based in part on observations at Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.